### v1

# Considering momentum dependence of heavy quarks interaction rate in calculating drag coefficient, and its impact on energy loss of b and c quarks

Abstract
This paper examines the effect of heavy particle momentum on their scattering frequency and calculates the drag coefficient by expanding the energy loss terms as polynomial expansions of momentum ($p$). This approach allows for a more precise investigation of momentum dependence of drag coefficient by incorporating the linear terms of these expansions. Additionally, the influence of particle’s momentum on radiative and collisional energy loss is more clearly determined. The study focuses on calculation of the nuclear modification factor ($R_{AA}$) of charm and bottom quarks in Pb-Pb collisions at $\sqrt{S_{NN}}=5.02\>TeV$. The results are compared with the latest data from ALICE and ATLAS experiments, conducted in 2021 and 2022. In this article, the initial distribution functions have been evolved numerically based on the Fokker-Planck equation.
pacs:
12.38.Bx, 12.39.-x, 14.65.Bt
I Introduction
The main goal of ultra-relativistic heavy ion collisions at the Relativistic Heavy-Ion Collider (RHIC) and the Large Hadron Collider (LHC) is to investigate QCD matter under extreme conditions. These experiments involve the collision of heavy ions, such as gold or lead, at very high energies to create a hot, dense environment of deconfined quarks and gluons, known as the Quark Gluon Plasma (QGP) STAR:2005gfr ; PHENIX:2004vcz ; QGP.Collins:1974ky . This highly excited state of matter, whose main constituents are light quarks and gluons, exhibits properties similar to a nearly perfect fluid. Therefore, it can be effectively described using hydrodynamic models while the system reaches local thermal equilibrium Ollitrault:2007du ; Becattini:2014rea .
Heavy quarks (such as charm and bottom) play an essential role in studying the properties of quark-gluon plasma as they are best probe for determining the medium’s transport properties HQ.vanHees:2005wb ; HQ.Rapp:2009my . These quarks are too heavy to be significantly produced through the interaction of thermal particles in the QGP. Therefore, the dominant mechanism for their production arises from hard interactions between partons within the colliding nuclei. Moreover, due to their large mass, these particles require a longer duration to reach thermal equilibrium with the environment. In some cases, they may even leave the plasma before achieving equilibrium. Thus, they are valuable probes to study the entire spacetime history of the deconfined medium. To gain a deeper insight into the properties of heavy quarks re-scattering in the QGP, one can analyze the evolution of their distribution function over time in the transverse momentum ($P_{T}$) plane using the Fokker-Planck (FP) equation.
Heavy quarks experience energy loss as they traverse within the hot, dense medium created by collisions. This energy dissipation occurs through two mechanisms: elastic collisions with constituents of the QGP and gluon bremsstrahlung or radiation resulting from interaction with other quarks, antiquarks, and gluons in the thermal environment. The interaction between heavy quarks and dynamic thermal media is characterized by transport coefficients. The drag coefficient, which quantifies the resistance experienced by a parton while moving through the plasma, is directly related to energy loss and can be calculated from the rate of energy loss per unit distance. Considerable efforts have been dedicated to achieving more accurate calculations of the drag and diffusion coefficients. One aspect of these researches examine the momentum dependence of the drag coefficient Mazumder:2011nj and momentum or temperature dependence of diffusion coefficient Hong:2023cwl ; Li:2019lex . Another part of these efforts is concentrated on developing various models for energy dissipation. There are several formulations available for calculating collisional Peshier:2006hi ; Braaten:1991jj ; Peigne:2008nd ; J. D. Bjorken ; Braaten:1991we and radiative energy loss Armesto:2003jh ; Armesto:2004vz ; Gyulassy:2000fs ; Djordjevic:2003zk . For a comprehensive review of many of these formalisms see Ref. Jamil:2010te .
To accurately calculate the drag coefficient, it is essential to consider the number of interactions that heavy particles undergo during propagation and the amount of energy lost in each interaction. As particles propagate through the thermal bath, their energy and momentum decrease, leading to a reduction in the number of both elastic and inelastic collisions over time. In this article, we determine the drag coefficient by considering energy loss coefficients as a linear function of particle’s momentum. This approach examines the impact of particle’s momentum on its collision rate, allowing for a more precise analysis of collisional and radiative energy losses. In this way, we analyze the drag coefficient as a function of momentum, addressing limitations in previous studies that overlooked or simplified its momentum dependence and lost some details.
This paper is organized as follows: In the next section, we provide a brief overview of the Fokker-Planck equation. Additionally, we discuss the space-time evolution of the system and outline the initial conditions. In Section (III), the drag and diffusion coefficients for collisional and radiative processes have been discussed. Our results are presented in Section (IV), and section (V) is dedicated to summary and discussions.
II System evolution
As mentioned before, heavy quarks reach thermodynamic equilibrium more slowly compared to lighter particles. Consequently, the interaction between heavy quarks and medium can be considered as interaction between equilibrium and non-equilibrium states. The equilibrium component, known as Quark-Gluon Plasma, is presumed to form at an initial temperature ($T_{i}$) and time ($\tau_{i}$) after the nuclear collision. The QGP expands and cools due to its high internal pressure, until transition to the hadronic phase at temperature $T_{c}$. Heavy quarks as the non-equilibrium component, exhibit Brownian motion within the thermal environment of the QGP Moore:2004tg ; Alam:1994sc . Therefore, Fokker-Planck equation, which is a simplified version of the Boltzmann equation for small momentum transformations Svetitsky:1987gq ; Rapp:2009my , provides a suitable framework for investigating the evolution of the heavy quark’s momentum distribution. According to this equation we have:
$\frac{\partial}{\partial t}f(p,t)=-\frac{\partial}{\partial p_{i}}\left[A_{i}(p)f(p,t)\right]+\frac{\partial^{2}}{\partial p_{i}\partial p_{j}}\left[D_{ij}(p)f(p,t)\right].$ (1)
Where $f(p,t)$ stands for the distribution function of heavy quarks and $A_{i}(p)$ and $D_{ij}(p)$ are related to the drag and diffusion coefficients, respectively. These coefficients characterize the interaction of heavy quarks with the QGP and are generally influenced by both temperature and momentum. The i and j indices represent different spatial directions, which can be omitted as we are considering a spatially uniform quark-gluon plasma in transverse plane. In the next section, we will discuss these coefficients in detail.
The initial transverse momentum distribution of charm and bottom quarks has been taken from Refs. Sheibani:2021ovo ; Modarres:2021gva ; Olanj:2020lkt . In Sheibani:2021ovo , the distribution functions are obtained using the APFEL legacy framework, incorporating the nCTEQ15 model and by considering the EMC effect to determine bound quarks within the nuclei at a specific center-of-mass energy. Subsequently, the Pythia8 simulation code is employed to calculate the initial spectra of heavy quarks after the Pb-Pb collision. The process of converting distribution functions from x-variable to transverse-momentum-dependent PDFs can be found in relation (1) of this article Sheibani:2021ovo .
As we mentioned, the QGP environment is constantly changing and evolving, and its temperature decreases during expansion. The temperature evolution of the background (QGP) is governed by relativistic hydrodynamics. Here, we consider the evolution equation of the QGP temperature, using the Bjorken flow in the Milne coordinate, as follows Chattopadhyay:2018apf ; Grozdanov:2015kqa :
$T(\tau)=T_{0}\left(\frac{\tau_{0}}{\tau}\right)^{\frac{1}{3}}\left[1+\frac{2}{3\tau_{0}T_{0}}\frac{\eta}{s}\left(1-\left(\frac{\tau_{0}}{\tau}\right)^{\frac{2}{3}}\right)\right]$ (2)
[…]
Furthermore, utilizing a temperature-dependent function for the running coupling $\alpha_{s}(T)$ is crucial, as temperature significantly influences the QCD coupling in the QGP system. We define $\alpha_{s}(T)$ as Braaten:1989kk :
$\alpha_{s}(T)=\frac{6\pi}{(33-2N_{f})\ln\left(\frac{19\>T}{\Lambda_{MS}}\right)}$ (3)
[…]
III Drag and diffusion coefficients
To solve the Fokker-Planck equation and derive the distribution function of heavy quarks at the transition temperature ($T_{c}$), we require the drag and diffusion coefficients. The diffusion coefficient can be calculated utilizing Einstein’s relation under conditions of weak coupling between the heavy quarks and the thermal bath Diffusion.Srivastava:2016igg ; Diffusion.Das:2010tj ; Diffusion.Akamatsu:2008ge . According to this relation, we have:
$D(p)=TA(p)E$ (4)
T is the temperature of the thermal bath and E is the energy of heavy quarks. Therefore, by calculating the drag coefficient, we will have both the drag and diffusion coefficients.
It is important to note that the drag coefficient carries information about the dynamics of heavy quark interactions with the medium and is influenced by the properties of the thermal bath. Consequently, the key aspect of determining the temporal evolution of heavy quark distribution function lies in calculating the drag force acting on them, or equivalently, the rate of energy loss per unit distance traveled by the heavy quark in the medium. So, we have:
$A(p)\propto-\frac{1}{p}\frac{dE}{dL}$ (5)
We have considered both kind of energy loss in calculating the drag coefficient. Therefore, it can be determined through the following relation:
$A(p)=-\frac{1}{p}\>\left[k_{1}\>(\frac{dE}{dL})_{coll}+k_{2}\>(\frac{dE}{dL})_{rad}\right]$ (6)
$k_{1}$ and $k_{2}$ are typically treated as constants, which can be derived from the best fit to the experimental results. These constants account for the number of collisions, degeneracy, and other factors.
However, to precisely determine the drag coefficient, it is necessary to consider the impact of particles’ momentum on the rate of their collisions, which results in energy loss. As the kinetic energy and momentum of heavy quarks decrease during propagation in QGP, the frequency of their collisions with plasma constituents is expected to decrease as well. Therefore, it can be said that the number of elastic and inelastic interactions, depends on the momentum of the particles.
Consequently, in this article, we have formulated the drag coefficient as follows:
$A(p)=-\frac{1}{p}\left[\>k^{\prime}(p)\left(\frac{dE}{dL}\right)_{coll}+\>k^{\prime\prime}(p)\left(\frac{dE}{dL}\right)_{rad}\>\right]$ (7)
The functions $k^{\prime}(p)$ and $k^{\prime\prime}(p)$ represent collisional and radiative energy loss coefficients, respectively. We assume these functions comprised of a constant value and a modified value which is expressed as polynomial expansions of momentum, as follows:
[…]
We consider these expansions up to first-order which include linear terms. By considering the energy loss coefficients as a function of $p$, the impact of particle’s momentum on the number of its interactions leading to collisional (elastic) and radiative (inelastic) energy losses can be examined separately. Setting $\alpha$ and $\beta$ to zero yields the initial expression of the drag coefficient in the equation (6).
Finally, by substituting (8) into (7), we have:
$A(p)=-\frac{1}{p}\left[\>k_{1}(1+\alpha\;p)\left(\frac{dE}{dL}\right)_{coll}+\>k_{2}(1+\beta\>p)\left(\frac{dE}{dL}\right)_{rad}\>\right]$ (9)
All coefficients including $k_{1}$, $k_{2}$, $\alpha$ and $\beta$ are determined through the fitting of experimental data. Figure (1) presents the drag coefficient for both cases of momentum-dependent coefficients and constant coefficients, at $T=250\;MeV$. As shown in the figure, incorporating the second expansion terms in the drag coefficient calculation results in a significant increase in its value at higher momenta. Consequently, we anticipate greater energy suppression in these regions. By comparing the $R_{AA}$ results obtained from above equation with the case where the energy loss coefficients are considered as constant values (expanded to the first term), the degrees of variation or improvement in the results can be measured.
In order to calculate collisional and radiative energy loss, a variety of methods have been introduced in recent studies. In our previous work Nezhad:2023mvb , we evaluated and compared three different approaches for calculating collisional energy loss, and found that the approach proposed by Braaten and Thoma, which utilizes the Hard-Thermal-Loop (HTL) framework Braaten:1991we , aligns better with recent experimental data. Therefore, we have employed the same method for calculating collisional energy loss in this work. Radiative energy loss has been calculated using the reaction operator formalism (DGLV) Gyulassy:2000fs ; Djordjevic:2003zk ; Wicks:2005gt and the generalized dead cone approach Saraswat:2015yll ; Saraswat:2017vuy .
Figure 1: Drag coefficient with momentum-dependent coefficients and constant coefficients at T=250 MeV.
IV Results
The nuclear suppression factor, $R_{AA}$, defined as the ratio of quenched to unquenched $P_{T}$ distribution, is derived from experimental measurements. This quenching phenomenon results from the interaction between heavy quarks and medium. This factor is influenced by several key processes, including the initial production of heavy quarks, the evolution of the medium, hadronization, and the interactions of heavy quarks within quark-gluon plasmas. We concentrate on the momentum-dependent drag coefficient and energy loss effects in QGP and estimate $R_{AA}$ as the ratio of the final heavy quark distribution in heavy-ion collisions to the initial heavy quark distribution in proton-proton collisions:
$R_{AA}(p_{T})\approx\frac{f_{f}(P_{T})^{A-A}}{f_{i}(P_{T})^{P-P}}$ (10)
The proportionality sign can be replaced with an equal sign by considering the ratio of the number of collisions and also degrees of degeneracy. We will adjust this coefficient by optimizing procedure. In this section, we calculate the nuclear suppression factor of charm and bottom quarks in a Pb-Pb collision at a center-of-mass energy of 5.02 TeV and compare our findings with recent LHC experimental data collected in 2021 and 2022. The parton distribution functions are evaluated using the Fokker-Planck equation, which we solve numerically at third-order relativistic hydrodynamics until the fireball reaches it’s freeze-out temperature of $T_{c}=155$ MeV. We calculate drag coefficient for both cases of first-order momentum expansion (9) and as a constant term (6). This procedure allows us to compare the results with experimental data and investigate the impact of momentum dependence of energy loss coefficients on results.
The final results are optimized to align with experimental data by adjusting free parameters like $\alpha$ and $\beta$ and minimizing the unweighted Chi-squared value:
$\chi^{2}=\sum_{i}\frac{\left[R_{AA}^{exp}(p_{T}(i))-R_{AA}^{th}(p_{T}(i))\right]^{2}}{\sigma_{i}^{2}}$ (11)
$R_{AA}^{exp}$ and $R_{AA}^{th}$ represent experimental and theoretical predictions for the suppression factor, respectively, while $\sigma$ indicates experimental error.
In our parameter optimization process, we utilize the Minuit package Minuit.James:1975dr , which is an effective and powerful tool that enables us to fit more free parameters and achieve high accuracy in minimizing the chi-squared value. This resource is essential for ensuring that our optimization efforts are both precise and reliable, thus enhancing the overall quality of our results.
Figure 2: Nuclear modification factor of charm quark in the (20–40%) centrality Pb-Pb collisions at $\sqrt{S_{NN}}=5.02$ TeV compared with ALICE 2021 data at forward rapidity ($42.5<y<$ 4) ALICE:2020sjb . The solid line represents the results obtained from the momentum-dependent coefficients, and dashed line corresponds to the drag coefficient with constant coefficients.
Figure 3: Nuclear modification factor of charm quark in the (30–50%) centrality Pb-Pb collisions at $\sqrt{S_{NN}}=5.02$ TeV compared with ALICE 2022 data at mid-rapidity ($|y|<0.5$) ALICE:2021rxa . The solid line represents the results obtained from the momentum-dependent coefficients, while the dashed line corresponds to the drag coefficient with constant coefficients.
Figures (2) and (3) illustrate $R_{AA}$ results for charm quark, derived from fitting theoretical predictions to the ALICE 2021 ALICE:2020sjb and ALICE 2022 data ALICE:2021rxa , respectively. Analysis of both figures indicates that the plot derived from the momentum-dependent drag coefficient, exhibits a more precise alignment with the experimental data. Furthermore, the $\chi^{2}$ value decreases from 0.3 to 0.2 for ALICE 2022 dataset and decreases from 0.47 to 0.44 for ALICE 2021 dataset, when we use momentum dependence coefficients.
The coefficients $\alpha$ and $\beta$ in Equation (9) for the c quark are found to be $\alpha=0.02$ and $\beta=0.05$. These values indicate that at higher momentum, both collisional and radiative energy losses of particles increase. Specifically, the impact of particle’s momentum on radiative energy loss becomes more significant and should be encountered. We also obtain $k_{1}=1.5$ and $k_{2}=2$ through the optimization process, which implies that the radiative energy loss for charm quarks is consistently greater.
Figure (4) shows our result for b quark fitted on ATLAS 2022 data ATLAS:2021xtw . Due to the limited amount of experimental data for the b quark, it is not feasible to accurately fit a model with too many free parameters. Therefore, in calculating drag coefficient for b quark, we have only considered the first term of the expansion with constant coefficients. By comparing theoretical results with the experimental data, the constant coefficients in (6) are found to be $k_{1}=5.5$ and $k_{2}=1.8$ Therefore, it seems that the collisional energy loss for b quark, is significantly higher than the radiative energy loss. We also have $\chi^{2}=0.124$ and as illustrated in the figure, the theoretical results exhibit a very good agreement with the experimental data.
As previously mentioned, examining higher-order terms in the drag coefficient is more challenging for b quark. However, by simulation and generating additional data point, the parameters $\alpha$ and $\beta$ for this quark are found to be similar to those of c quark. The difference is that for the b quark, even by considering higher order coefficients, we find $k_{1}$ greater than $k_{2}$. The values of $k_{1}$ are found to lie within the approximate range of $32.5<k_{1}<$ 3 and $k_{2}$ is in the approximate range of $1.81<k_{2}<$ 1.8 . These values indicates that for the bottom quark, collisional energy loss is dominant mechanism for both calculation of the drag coefficient. This is mainly due to its large mass which suppresses gluon radiation Dokshitzer:2001zm .
Figure 4: Nuclear modification factor of bottom quark in the (40–60%) centrality Pb-Pb collisions at $\sqrt{S_{NN}}=5.02$ TeV compared with ATLAS 2022 data ATLAS:2021xtw . This graph is obtained from considering the drag coefficient with constant coefficients.
V Summary and discussions
In this paper, we present a comprehensive analysis of the drag coefficient within the framework of the Fokker-Planck equation. We account for the momentum dependence of heavy quarks scattering rate when calculating the drag coefficient. This is achieved by modeling the energy loss coefficients as a function of momentum. The results are then compared with a simpler form of drag coefficient which is linked to energy loss through constant coefficients. Our findings indicate that by incorporating linear momentum-dependent terms in the drag calculation, we can more accurately identify the radiative and collisional energy losses for particles with varying momenta, and $R_{AA}$ results demonstrate a better agreement with experimental data. The findings show that as momentum increases, the energy loss also rises, and this enhancement is more significant in radiative energy loss. Additionally, our analysis of $b$ and $c$ quarks reveals that for the $c$ quark, radiative energy loss exceeds collisional energy loss. While, for the $b$ quark, due to its higher mass, collisional energy loss is more prominent than radiative energy loss. Nevertheless, accurately determining free parameter of our model in evolution of $b$ quark is challenging due to the limited amount of experimental data. With more experiments in the future and an expanded dataset for the b quark, we will be able to accurately determine the $\alpha$ and $\beta$ coefficients. This will enable us to quantify the influence of particle’s momentum on scattering rates and, consequently, the associated energy loss.